OpenStudy (anonymous):

Is sqrt(a/b) = to (sqrt a)/(sqrt b) = sqrt a * (1/(sqrt b))?

OpenStudy (anonymous):

\[\sqrt{\frac{ a }{ b }}= \frac{ \sqrt{a} }{ \sqrt{b} }= \sqrt{a}*\frac{ 1 }{ \sqrt{b} }\]

OpenStudy (anonymous):

This is true right?

OpenStudy (campbell_st):

well the solution you have is similar to \[\frac{3}{4} = 3 \times \frac{1}{4}\]

OpenStudy (campbell_st):

and you have correctly applied the rule for dividing radicals

OpenStudy (campbell_st):

so it all makes sense to me

OpenStudy (campbell_st):

but you need the proviso... \[b \ne 0\]

OpenStudy (anonymous):

yes thank you

OpenStudy (zarkon):

\[a\ge0,b>0\]

OpenStudy (anonymous):

oh yea cuz b and a have to be positive